if i sold brownies for 3.00 and cookies for 2.50 i sold 3 less cookies and the total is 218.00. ive been working on this for an hour along with my parents. we just want the equation.
Answer by richwmiller(17219)

b=c+3
b*3+c*2.5=218
b=41 c=38
check
3*41+2.5*38=218
ok